OpenStudy (anonymous):

how do you right this in slope-intercept form x - 2y = 14

OpenStudy (texaschic101):

slope intercept form : y = mx + b x - 2y = 14 (subtract x from both sides) -2y = -x + 14 (multiply the equation by -1 to make y positive) 2y = x - 14 (now divide by 2) y = 1/2x - 7 <-- slope intercept form any questions ?

OpenStudy (anonymous):

thanx

OpenStudy (texaschic101):

anytime...glad to help :)